Al Simpson helped start Excel System in 2010. At the time, he purchased 119 000 shares of stock at $1 per share. In 2015. he has the opportunity to sell his interest in the Folsom Corp. for $50 a share in cash. His capital gain tax would be 10 percent.
a. If he sells his interest, what woll be the value for before profit, taxes, and aftertax?
b. Assume instead of cash. he accept Folsom Corp. stock valued st $50 per share. he pays tax at that time. he hold the stock for five year and sell it for $83.50. the stock pays no cash dividends. what will the value for before tax profit, taxes. and aftertax be in 2020. tax gains 10 percent.
c. what is the percent value year 2015. with 9 percent discount.
Answer #1

a )   119000 Shares purchased at $ 1 per share in 2010. later in 2015 has opportunity to sell shares for $ 50

Calculation of capital gain

Sale consideration (119000*50) =         5950000

Initial cost                              =            119000

Profit before tax                      =        5831000

Tax (10%)                     =             583100

Profit after tax                       =          5712000

b) Assume instead of cash. he accept Folsom Corp. stock valued st $50 per share. he pays tax at that time. he hold the stock for five year and sell it for $83.50.

Calculation of gain

Sale consideration (119000*83.50) = 9936500

Initial investment =                             5950000

Profit before tax =                              3986500

Tax ( 10%)                                          398650

Profit after tax                                   3587850

c) Percent value in 2015 with discount rate 9%

Initial investment in 2010 = 119000

Value in 2015 = 119000 PVIF ( 9%,5) = 183096

percent value = (183096/119000)*100

                        153.86 %